# ConsistTalk: Intensity Controllable Temporally Consistent Talking Head Generation with Diffusion Noise Search

### v1

@@ Abstract
Introduction
Audio-driven portrait animation aims to synthesize expressive talking head videos from a static portrait image and accompanying speech audio. Early works (sadtalker; aniportrait) primarily employ a two-stage pipeline using 3D motion or keypoints as intermediate representations. In the first stage, audio features are extracted and mapped to motion latents, which are then used to generate video frames through a rendering network. However, representational bottlenecks limit the ability of these methods to produce dynamic and realistic facial movements.
@@ Introduction
Methodology
In this section, we present ConsistTalk for audio-driven portrait image animation, designed to mitigate flickering and temporal inconsistency problems effectively. Let $\boldsymbol{I}_{ref}$ denote the reference image, and let $\{\boldsymbol{I}_{t-1},\cdots,\boldsymbol{I}_{t-n}\}$ represent the preceding $n$ generated frames. Our architecture is built upon the Dual-UNet design (emo; hallo; echomimic). We first use CLIP (clip) to extract features from the reference ID portrait image. Compared with Hallo(hallo)’s face embedding, it is compatible with the stable diffusion’s text-image cross-attention space. To decouple temporal motion dynamics from identity information in motion frames, we introduce a facial optical flow-guided temporal module (Sec. Optical Flow Guided Temporal Module). In addition, a multimodal joint modeling module is proposed to learn audio-video intensity from both audio signals and facial/head motion sequences (Sec. Motion Intensity Sequence Modeling). The intensity value is then used to guide noise initialization, enhancing the quality of generation and ensuring temporal consistency (Sec. Intensity Guided Diffusion Noise Search).
Optical Flow Guided Temporal Module
@@
Motion Intensity Sequence Modeling
@@
Intensity Guided Diffusion Noise Search
@@
$R(\boldsymbol{z}_{l},\boldsymbol{z}_{ref,\tau},\boldsymbol{z}_{l-1,\tau})=$
$\lambda_{low}\cdot\left\langle\boldsymbol{z}_{l}^{low},\boldsymbol{z}_{ref,\tau}^{low}\right\rangle-\lambda_{high}\cdot\left\langle\boldsymbol{z}_{l}^{high},\boldsymbol{z}_{l-1,\tau}^{high}\right\rangle$ (2)
@@
Figure 4: Qualitative comparisons with State-of-the-Art talking head generation methods.
Experiment
Experimental Setups
@@
We compare our method against state-of-the-art audio-driven talking head approaches with publicly available implementations. For non-diffusion models, we compare with SadTalker (sadtalker). And for diffusion-based models, we compare with AniTalker (anitalker), Hallo(v1&v2) (hallo; hallo2), EchoMimic (echomimic), Sonic (sonic) and FantasyTalking (fantasytalking).
@@
Results and Analysis
Quantitative Comparison. We show the quantitative comparison results in Table 1. ConsistTalk consistently outperforms or matches state-of-the-art methods across multiple metrics. Regarding overall video quality, our framework achieves the lowest FVD score of 171.7, representing a 16.7% reduction compared to the second-best model, Sonic (206.1), demonstrating superior generation quality. Facial fidelity, measured by E-FID, also shows strong performance at 1.4397, narrowly trailing the best prior work. Our method also achieves the lowest ID-Dist score, underscoring its strength in identity preservation and validating the effectiveness of our IC-Init strategy. For temporal stability, as measured by Flicker (0.4218), our model reduces flicker by over 22% compared to the next best, SadTalker, and by nearly 39% compared to Sonic. Additionally, VBench scores of 99.68 and 98.28 further confirm superior cross-frame background steadiness. Lip synchronization quality remains high, with a Sync-C score of 6.83, approximately 13.2% higher than the baseline model Hallo (6.3), indicating that the A2I module preserves accurate mouth movements. Moreover, the model achieves the highest Beat Align (1.659) and Diversity (3.48) scores, reflecting more varied and better-synchronized head motions with audio.
@@ Conclusion
We leave additional results and analyses, videos, together with discussions considering limitations and ethical considerations in the supplementary materials.

### v2

@@ Abstract
1 Introduction
Audio-driven portrait animation aims to synthesize expressive talking head videos from a static portrait image and accompanying speech audio. Early works (sadtalker; aniportrait) primarily employ a two-stage pipeline using 3D motion or keypoints as intermediate representations. However, representational bottlenecks limit the ability of these methods to produce dynamic and realistic facial movements.
@@ 2 Related Work
3 Methodology
In this section, we present ConsistTalk for audio-driven portrait image animation, designed to mitigate flickering and temporal inconsistency problems effectively. Let $\boldsymbol{I}_{ref}$ denote the reference image, and let $\{\boldsymbol{I}_{t-1},\cdots,\boldsymbol{I}_{t-n}\}$ represent the preceding $n$ generated frames. Our architecture is built upon the Dual-UNet design (emo; hallo; echomimic). We first use CLIP (clip) to extract features from the reference ID portrait image. Compared with Hallo(hallo)’s face embedding, it is compatible with the stable diffusion’s text-image cross-attention space. To decouple temporal motion dynamics from identity information in motion frames, we introduce a facial optical flow-guided temporal module (Sec. 3.1). In addition, a multimodal joint modeling module is proposed to learn audio-video intensity from both audio signals and facial/head motion sequences (Sec. 3.2). The intensity value is then used to guide noise initialization, enhancing the quality of generation and ensuring temporal consistency (Sec. 3.3).
3.1 Optical Flow Guided Temporal Module
@@
3.2 Motion Intensity Sequence Modeling
@@
3.3 Intensity Guided Diffusion Noise Search
@@
$R(\boldsymbol{z}_{l,T-T^{\prime}},\boldsymbol{z}_{ref,\tau},\boldsymbol{z}_{l-1,\tau})=$
$\lambda_{low}\cdot\left\langle\boldsymbol{z}_{l,T-T^{\prime}}^{low},\boldsymbol{z}_{ref,\tau}^{low}\right\rangle-\lambda_{high}\cdot\left\langle\boldsymbol{z}_{l,T-T^{\prime}}^{high},\boldsymbol{z}_{l-1,\tau}^{high}\right\rangle$ (2)
@@
Figure 4: Qualitative comparisons with State-of-the-Art talking head generation methods on HDTF (hdtf) dataset.
4 Experiment
4.1 Experimental Setups
@@
We compare our method against state-of-the-art audio-driven talking head approaches with publicly available implementations: SadTalker (sadtalker), AniTalker (anitalker), Hallo(v1&v2) (hallo; hallo2), EchoMimic (echomimic), Sonic (sonic) and FantasyTalking (fantasytalking).
@@
4.2 Results and Analysis
Quantitative Comparison. We show the quantitative comparison results in Table 1. ConsistTalk outperforms or matches state-of-the-art methods across multiple metrics. Regarding overall video quality, our framework achieves the lowest FVD score of 171.7, representing a 16.7% reduction compared to the second-best model, Sonic (206.1), demonstrating superior generation quality. Our method also achieves the lowest ID-Dist score, underscoring its strength in identity preservation and validating the effectiveness of our IC-Init strategy. For temporal stability, as measured by Flicker (0.4218), our model reduces flicker by over 22% compared to the next best, SadTalker, and by nearly 39% compared to Sonic. Additionally, VBench scores of 99.68 and 98.28 further confirm superior cross-frame background steadiness. Lip synchronization quality remains high, with a Sync-C score of 6.83, approximately 13.2% higher than the baseline model Hallo (6.3), indicating that the A2I module preserves accurate mouth movements. Moreover, the model achieves the highest Beat Align (1.659) and Diversity (3.48) scores, reflecting more varied and better-synchronized head motions with audio.
@@ 5 Conclusion
Acknowledgements
This work was supported by National Natural Science Foundation of China 62376255.